# Rethinking Weight-Averaged Model-merging

### v2

@@ 5.2 Variance Reduction in Model Weight Averaging
$\small\text{Var}(\overline{\mathbf{W}})=\text{Var}\left(\frac{\mathbf{W}_{A}+\mathbf{W}_{B}}{2}\right)$ (13)
$=\frac{1}{4}\left[\text{Var}(\mathbf{W}_{A})+\text{Var}(\mathbf{W}_{B})\right]=\frac{1}{4}(\sigma_{A}^{2}+\sigma_{B}^{2})$
Mathematical Justification. We now compare the variance of the averaged weights $\overline{\mathbf{W}}$ with the variances of the individual models’ weights $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$. The variances for models A and B are $\sigma_{A}^{2}$ and $\sigma_{B}^{2}$, respectively. If $\sigma_{A}=\sigma_{B}=\sigma$, then $\text{Var}(\overline{\mathbf{W}})=\frac{1}{2}\sigma^{2}$. This shows that, when the variances of the two models are equal, the variance is reduced by a factor of two. More generally, since the average of the variances is always less than or equal to the maximum variance, we have:
$\small\frac{1}{4}(\sigma_{A}^{2}+\sigma_{B}^{2})\leq\max(\sigma_{A}^{2},\sigma_{B}^{2})$ (14)
Thus, the variance of the averaged model $\overline{\mathbf{W}}$ is always less than or equal to the larger of $\sigma_{A}^{2}$ and $\sigma_{B}^{2}$.
@@ 5.3 How Weight Magnitudes and Variance Effects Model Outputs
According to Wang et al. [32], we have the following property 1 and theorem 1 for the connections between model weights’ magnitudes/variances and outputs’ magnitudes/variances:
Property 1: Let $f_{\theta}(x)$ be a fully-connected neural network, where $\theta$ denotes the set of parameters. Assume the activation function $\phi$ is Lipschitz continuous with constant $L$, and the weight matrices $\mathbf{W}^{(m)}$ are random matrices with independent and identically distributed (i.i.d.) sub-Gaussian entries. Then, for any input $\mathbf{x}$, the output at the $m$-th layer, $\mathbf{y}^{(m)}$, satisfies the following upper bound:
@@
with a probability of at least $\left(1-2e^{-\tau^{2}}\right)^{m}$, where $(m)$ indexes the neural network layer, with m representing the exponential operation, and $\lambda=\sqrt{N}+C_{s}K_{s}^{2}(\sqrt{N}+\tau)$, with $N$ being the maximum number of neurons across all layers, $C_{s}$ representing a constant, and $K_{s}=\max\left\|\mathbf{W}^{(m)}\right\|_{2}$. This property shows that the output norm can grow exponentially with the number of layers $m$, controlled by the Lipschitz constant of the activation function and the bound on the weight matrices.
Theorem 1: Assume the weights in each layer are i.i.d. sub-Gaussian random variables with variance $\sigma_{w}^{2}$, and the biases are with variance $\sigma_{b}^{2}$. Furthermore, let the activation function $\phi$ be Lipschitz continuous with constant $L$. Then, the variance of the output of the final layer, represented by $V(f_{\theta}(x))$, can be bounded by:
$\small\begin{split}V(f_{\theta}(x))\leq&\|x\|_{2}^{2}(L^{2}N)^{m}\prod_{m=1}^{M_{l}}(\sigma^{(m)}_{w})^{2}+L^{2}N\left(\sigma_{b}^{(M_{l})}\right)^{2}+\\ &L^{2}\sum_{m=1}^{M-1}\left\{N\left(\sigma_{b}^{(m)}\right)^{2}\prod_{l=m+1}^{M}\left[L^{2}N\left(\sigma^{(m)}_{w}\right)^{2}\right]\right\},\end{split}$ (16)
where $M_{l}$ denotes the maximum number of layers, and $L$ represents the Lipschitz constant. This inequality shows that the output variance grows with the depth $M$ of the network and the variances of the weights and biases.
Property 1 and Theorem 1 show that the increase of model weights’ magnitudes/variances will also magnify outputs’ magnitudes/variances and vice-versa. As shown in Fig. 5, consider a neural network with input $\mathbf{x}$ and output $\mathbf{y}$, where weights are denoted as $\mathbf{W}$. When model weights have large magnitudes and high variance, predictions become unreliable. This issue can arise from several factors, including task specificity, lack of regularization or improper weight initialization. Together, these factors can drive the weights to unstable values, reducing the model’s ability to generalize effectively. On the other hand, when $\mathbf{W}\to 0$, the output $\mathbf{y}$ will tend to a fixed point determined predominantly by the network’s bias terms, thus becoming largely independent of the input $\mathbf{x}$. Hence, if the magnitude of $\mathbf{W}$ is small, but not identically zero, then the range of outputs for $\mathbf{y}$ will be limited. This reduces the sensitivity of $\mathbf{y}$ to changes in $\mathbf{x}$, effectively shrinking the output space (similar to the L1 or L2 regularization). Such behavior can enhance the robustness of the model, as smaller weight magnitudes reduce the model’s response to minor perturbations in input $\mathbf{x}$. However, this comes with a trade-off: the expressive capacity of the model diminishes as well, potentially leading to poor performance on complex tasks.
@@
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 3: As the aforementioned analysis and experimental study in Sec. 6.3, model-merging is more robust to the increase in weights’ magnitudes/variances than model ensemble, by bringing down the maximum magnitudes/variances of individual models before merging. This is because, generally, the small magnitudes of weights enhance robustness by minimizing the network’s sensitivity towards input variations, but at the cost of reducing model expressiveness. In this case, model merging is similar to regularization. This balance between robustness and model expressiveness is crucial and must be carefully tuned based on the specific task requirements. For certain tasks where noise resilience is more valuable than fine-grained feature sensitivity, lower weight magnitudes may improve performance. However, for tasks requiring detailed pattern recognition, lower weight magnitudes may hinder the network’s ability to capture critical pattern distinctions in the data.
@@ 6 Experiments
The experimental settings can be found in the supplementary materials.
@@ 6.1 The Patterns Contained in Model Weights
As shown in Fig. 1(a,b), the visualization of the linear classifier on Cifar100 clearly shows the patterns of the visual class learned, which are strongly correlated with the average of all images within each class. The template of each class acts as a merged collection of representative images of the respective class. Interestingly, if we merge the weights of the first 5 classes in Fig. 1 with the last 5 classes, it gives us the results of Fig. 2. As we analyzed before, the weight-averaged model-merging acts as the Mixup augmentation, which is a linear combination of different matrices, in the weight space. For instance, when the “bicycle” weight matrix is combined with the “bottle” weight matrix, we can see the wheels of bikes combined with a vertical bottle. We can also observe that after merging, each element within the combined template noticeably fades. This phenomenon can also explain the limitation of expressiveness of the merged model as described in Sec. 4.
Fig. 6 visualizes the linear matrices and training image averages for some Tiny ImageNet classes. For example, “goldfish” shows a reddish center in the averaged images and also shows a reddish center in the classifier weights; similarly, we can observe the eyes and nose of “koala”; for pretzel, we can even see the shape of the cracker from the weight visualization. However, they are not as clear as for Cifar100. This may be due to the relatively large dataset size for each class for this dataset, as the average of all images in a class itself is not very clear. This poor clarity can also be explained by the limited expressiveness of a single linear classifier and the more complicated patterns present in Tiny ImageNet than Cifar100. Besides the linear classifier visualization, we also conduct common deep learning models with nonlinear weights visualization by viewing the first layer of convolutional kernels (7 $\times$ 7 in size) from the ResNet50 model [8] in Fig. 7 (VGG model weight visualization can be found in the supplementary materials). We can see the patterns of texture and edges for the basic level geometric feature detection is similar to the Haar features described in Sec. 3.1.
The visualization for every class in Cifar100 and Tiny ImageNet is put in the supplementary materials.
@@ 6.2.1 Observations over Different Models
As shown in Fig. 3 for the merging of 10 models (the experiments for merging between 2 and 7 models are in the supplementary materials, where “Perf Ave” denotes the average performance of individual models; “deit_tiny” and “vit_tiny” denote deit_tiny_patch16_224 and vit_tiny_patch16_224 models, respectively), different structures have different performances, but model ensemble on logits generally perform the best across different models [8, 9, 22, 24, 6, 27]; while uniform soups perform the worst. We can also discover that greedy soups are generally better than uniform soups due to their mechanism to greedily select the best performing model to merge. It ensures that a good result can be reached by only incorporating the first model. The greedy ensemble on both logits and features also produces results that are comparable to greedy soups.
In general, adding more models into the merging process leads to better performance for all methods, including model-merging and ensemble. Interestingly, the ViT model, including 3 ViT-based architectures in the experiments, shows large gaps between uniformly averaging and logits ensemble, where such gaps suffer large increases by adding more models into the soups. This may happen because of the architectural differences between convolutional neural networks (e.g., DenseNet, ResNet and VGG) and Vision Transformers (ViTs). CNNs are more adapted to the learning of local patterns within different scales of images, making it ideal for model merging. On the contrary, ViTs have global attention mechanisms, where the templates inside models represent attention and correlations between tokens, which makes it hard for them to work for model merging. ViT models also show similar results on other data (as shown by the DeiT on TinyImageNet results, which can be found in the supplementary materials).
@@
We also visualize the gaps between Logits Ensemble and Uniform Soups Across Different Configurations (as a function of the number of merged models and different model architectures, but excluding ViT models as their gaps are too large) in Fig. 8. We can see that the more models in the merging process, the larger the gaps between logits ensemble and uniform soups. The gaps are different according to different models, but generally, there is a downtrend when models contain more layers (e.g., ResNet18 to ResNet152, DenseNet121 to DenseNet201, VGG11 to VGG19 without BatchNorm). This may have been caused by the trend of performance saturation when more layers are included, as the model will be empowered with stronger fitting abilities. Batch normalization can break this downtrend. We have also examined the gaps between logits ensemble and greedy soups (in the supplementary materials), which shows that greedy soups can alleviate this marginal effect.
@@ 6.2.2 Observations over Different Datasets
For the dataset-wise examination, we perform model merging and ensemble on 7 prevalent datasets with different data sizes. From Tab. 1, we can observe ensemble still generally performs better than merging. Different datasets have different performances and this is due to the characteristics of each dataset. Similarly to the model-wise comparison, with more models in the merging (2-7 models merging results in the supplementary materials), the better the performance for both merging and ensembling. The model ensembling is more stable than model merging in terms of the best and worst performance on different data. For example, on the PathMNIST [37, 38] dataset, uniform soup merging does not work really well compared to the relative performance on the Cifar10/Cifar100 or Tiny ImageNet datasets (that only achieves 40.85% accuracy, which is largely behind model ensemble methods). A similar situation happens on Celeba [15] dataset: the uniformly average model only achieves 53.19%. Nevertheless, the greedy soup models can mitigate this problem (88.48% on PathMNIST and 93.06 on Celeba) as they can ensure a decent result by only considering a single model. These results may be caused by the contradictions between the two models’ weights.
@@ 6.2.3 Merging Models trained over different Tasks
Table 2: Performance Comparison for Task Merging Across Different Datasets with Various Model Averaging Methods. Model 1 is the VGG19 model trained on the Cifar100 dataset, while Model 2 denotes the VGG19 model trained on each different dataset with different tasks. Feature ensemble is better than logits ensemble.
Datasets Size Model1 Model2 Uniform Soups Ens Lgt Ens Features
Cifar10 60,000 70.37 5.84 49.63 69.93 69.93
PathMNIST 107,180 70.37 1.29 24.25 69.88 70.00
DermaMNIST 10,015 70.37 5.77 60.37 69.06 69.36
Celeba 202,599 70.37 1.03 49.33 69.48 69.61
TinyImageNet 100,000 70.37 18.16 61.02 58.40 58.52
ChestXRay 112,120 70.37 15.18 61.38 68.18 68.33
We also explore model merging between different well-trained models on different tasks, as shown in Tab. 2. Generally, merging within Tiny ImageNet (where Model 2 was trained on) has the best performance of 61.02% accuracy, which is better than the ensemble on logits (58.40% accuracy). This is because the Tiny ImageNet dataset is a general and relatively large computer vision dataset, containing a diverse set of images. Interestingly, Cifar10 should be the most similar one to Cifar100, but the merging results do not perform better than the model ensemble, as it is too small. On the contrary, ChestXRay dataset should be very different from Cifar100, but it has a reasonably decent improvement. This may illustrate that the dataset size also matters and the potential high similarity of convolutional kernels on both data was learned. However, the performance on Celeba is not good. This may be caused by the Celeba dataset only containing human faces, which can be seen as a fine-grained classification task. Due to the uniformity of the data format within Celeba, it contradicts the diverse and rich general computer vision features learned by the model weights on Cifar100.

### v3

@@ 5.2 Weight Variance Reduces while Averaging
$\small\text{Var}(\overline{\mathbf{W}})=\text{Var}\left(\frac{\mathbf{W}_{1}+\mathbf{W}_{2}}{2}\right)$ (13)
$=\frac{1}{4}\left[\text{Var}(\mathbf{W}_{1})+\text{Var}(\mathbf{W}_{2})\right]=\frac{1}{4}(\sigma_{1}^{2}+\sigma_{2}^{2})$
Mathematical Justification. We now compare the variance of $\overline{\mathbf{W}}$ to the variances of $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$. If $\sigma_{1}=\sigma_{2}=\sigma$, then $\text{Var}(\overline{\mathbf{W}})=\frac{1}{2}\sigma^{2}$. This shows that, when the variances of the two models are the same, the variance is reduced by a factor of two. More generally, we have:
$\small\frac{1}{4}(\sigma_{1}^{2}+\sigma_{2}^{2})\leq\max(\sigma_{1}^{2},\sigma_{2}^{2})$ (14)
Thus, the variance of the averaged model $\overline{\mathbf{W}}$ is always less than or equal to the larger member between $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$.
@@ 5.3 How Weight Magnitudes and Variance Effects Model Outputs
We have the following property 1 and theorem 1 to build the connections between model weights’ magnitudes/variances and outputs’ magnitudes/variances, following  [34]:
Property 1: Let $f_{\theta}(x)$ be a fully-connected neural network, where $\theta$ is the parameters. Assume the activation function $\phi$ is Lipschitz continuous with constant $L$, and the weight matrices $\mathbf{W}^{(m)}$ are random matrices with independent and identically distributed (i.i.d.) sub-Gaussian entries. Then, for any input $\mathbf{x}$, the output at the $m$-th layer, $\mathbf{y}^{(m)}$, satisfies the following upper bound:
@@
with a probability of at least $\left(1-2e^{-\tau^{2}}\right)^{m}$, where $(m)$ indexes the neural network layer, with m representing the exponential operation, and $\lambda=\sqrt{N}+C_{s}K_{s}^{2}(\sqrt{N}+\tau)$, with $N$ being the maximum number of neurons across all layers, $C_{s}$ representing a constant, and $K_{s}=\max\left\|\mathbf{W}^{(m)}\right\|_{2}$. This property shows that the output norm can grow exponentially with the number of layers $m$, controlled by the Lipschitz constant of the activation function and bounded by the max entry value of weight matrices.
Theorem 1: Assume the weights in each layer are i.i.d. sub-Gaussian random variables with variance $\sigma_{w}^{2}$, and the biases are with variance $\sigma_{b}^{2}$. Also, let the activation function $\phi$ be Lipschitz continuous with constant $L$. Then, the variance of the output of the final layer, represented by $\mathbb{V}(f_{\theta}(x))$, can be bounded by:
$\small\begin{split}\mathbb{V}(f_{\theta}(x))\leq&\|x\|_{2}^{2}(L^{2}N)^{m}\prod_{m=1}^{M_{l}}(\sigma^{(m)}_{w})^{2}+L^{2}N\left(\sigma_{b}^{(M_{l})}\right)^{2}+\\ &L^{2}\sum_{m=1}^{M-1}\left\{N\left(\sigma_{b}^{(m)}\right)^{2}\prod_{l=m+1}^{M}\left[L^{2}N\left(\sigma^{(m)}_{w}\right)^{2}\right]\right\},\end{split}$ (16)
where $M_{l}$ denotes the maximum number of layers, and $L$ represents the Lipschitz constant. This inequality shows that the output variance grows with the depth of the network, as well as impacted by the variances of the weights / bias.
Property 1 and Theorem 1 show that the increment of model weights’ magnitudes/variances will also magnify outputs’ magnitudes/variances. As shown in Fig. 5, consider a neural network with input $\mathbf{x}$ and output $\mathbf{y}$, where weights are denoted as $\mathbf{W}$. When model weights have large magnitudes and high variance, predictions become unreliable. This issue can arise from several factors, including task specificity, lack of regularization or improper weight initialization. Together, these factors can drive the weights to unstable values, reducing the model’s ability to generalize effectively. On the other hand, when $\mathbf{W}\to 0$, the output $\mathbf{y}$ will tend to a fixed point determined predominantly by the network’s bias terms, becoming largely independent of the input $\mathbf{x}$. Hence, if the magnitude of $\mathbf{W}$ is small, but not identically zero, then the range of outputs for $\mathbf{y}$ will be constrained. This reduces the sensitivity of $\mathbf{y}$ to the changes of $\mathbf{x}$, effectively shrinking the output space. In this case, model-merging acts similar to the L1 or L2 regularization. Such behavior can enhance the robustness of the model, as smaller weight magnitudes reduce the model’s response to minor perturbations in input $\mathbf{x}$. But this comes with a trade-off: the expressive capacity of models may diminish as well, potentially leading to compromised performance on complex tasks.
@@
[backgroundcolor=white!10,rightline=true,leftline=true,topline=true,bottomline=true,roundcorner=2mm,everyline=true,nobreak=false] Proposal 3: As the aforementioned analyses and experimental study in Sec. 6.3, model-merging is more robust to the changes in weights’ magnitudes/variances than model ensemble, by bringing down the maximum magnitudes/variances of individual models before merging. This is because, generally, the small magnitudes of weights enhances robustness by minimizing the network’s sensitivity towards input variations, but somehow reduces the expressiveness from ensembling all models. From this perspective, model-merging acts similarly as regularization. This balance between robustness and expressiveness is crucial and should be carefully considered based on the specific task requirements. For certain tasks where noise resilience is more valuable than fine-grained feature sensitivity, lower weight magnitudes can improve performance.
@@ 6 Experiments
The experimental settings and details can be found in the supplementary materials.
@@ 6.1 The Patterns Contained in Model Weights
As shown in Fig. 1(a,b), the visualization of the linear classifier on Cifar100 explicitly shows the patterns per classifier learned, which are strongly correlated with the average of images per class. The patterns learned per class acts as a fused representative image. Interestingly, if we merge the weights of the first 5 classes with the last 5 classes in Fig. 1, it gives us the results of Fig. 2. As mentioned, the weight-averaged model-merging acts as a Mixup operation, which is a linear combination of different matrices, in the weight space. For instance, when the “bicycle” weight matrix is combined with the “bottle” weight matrix, we can see the wheels of bikes combined with a vertical bottle. We can also observe that, after merging, each element within the combined template noticeably fades. This phenomenon can also explain the limitation of expressiveness of the merged model as described in Sec. 4.
Fig. 6 visualizes the image average per class and its corresponding weight for Tiny ImageNet classes. For instance, “goldfish” shows a reddish center in the averaged images and also shows a reddish center in the classifier weights; similarly, we can observe the eyes and nose of “koala”; for pretzel, we can see the shape of the cracker from its weight visualization. However, they are not as clear as for Cifar100. This may be due to the relatively large dataset size, as the average of all images in a class itself is also vaguer. This compromised clarity can also be caused by the limited expressiveness of single linear classifiers and there are more complicated patterns present in Tiny ImageNet than Cifar100. Besides the linear classifier visualization, we also conduct common deep learning models with nonlinear weights visualization by viewing the first layer of convolutional kernels (7 $\times$ 7 in size) from the ResNet50 model [9] in Fig. 7 (VGG model weight visualization can be found in the supplementary materials). We can see the patterns of texture and edges for the basic level geometric feature detection is similar to the Haar features described in Sec. 3.1.
The visualizations for all classes within Cifar100 and Tiny ImageNet are put in the supplementary materials.
@@ 6.2.1 Observations over Different Models
As shown in Fig. 3 for the merging of 10 models (the experiments for merging from 2 to 7 models are in the supplementary materials). For abbreviations, “Perf Ave” denotes the average performance of individual models; “deit_tiny” and “vit_tiny” denote deit_tiny_patch16_224 and vit_tiny_patch16_224 models, respectively. Different models show different merging/ensembling performances, but model ensemble on logits generally performs the best across different tested architectures [9, 10, 24, 26, 5, 29]. However, uniform soups perform the worst. We can also discover that greedy soups generally perform better than uniform soups due to their mechanism to greedily select the best performing model to merge. In fact, it guarantees that reasonably promising results can be achieved by only incorporating the first model (if the later-on merged model performance is not ideal). The greedy ensemble on both logits and features produces results that are comparable to greedy soups.
In general, adding more models into the merging process leads to better performance for all methods, including model-merging and ensembling. Interestingly, the ViT model, including 3 ViT-based architectures in the experiments, shows large gaps between uniformly averaging and logits ensembling, where such gaps suffer larger increases by adding more models into. This happens may be caused by the architectural differences between convolutional neural networks and Vision Transformers (ViTs). CNNs are more adapted to the learning of local patterns within different scales of images, making it ideal for merging. On the contrary, ViTs have global attention mechanisms, where the templates represent attention and correlations between tokens, making them sensitive for small changes. ViT models also show similar deterioration on other data (as shown by the DeiT on TinyImageNet and it can be found in the supplementary materials).
@@
We also visualize the gaps between Logits Ensemble and Uniform Soups Across Different Configurations (excluding ViT models as the gaps are too large, so in different scales) in Fig. 8(a). We can see that the more models in the merging process, the larger the gap between logits ensemble and uniform soups. The gaps are different according to different models, but generally, there is a downtrend when models contain more layers (e.g., ResNet18 to ResNet152, DenseNet121 to DenseNet201, VGG11 to VGG19 without BatchNorm). This may be caused by the performance saturation when more layers are included, as the model will be empowered with stronger fitting abilities. But batch normalization can break this trend. We have also examined the gaps between logits ensemble and greedy soups in the supplementary materials, showing that greedy soups can alleviate this marginal effect.
@@ 6.2.2 Observations over Different Datasets
For the dataset-wise examination, we perform model-merging and ensembling on 7 prevalent datasets with different data sizes (TinyImageNet and ChestXRay are relatively large, and both with more than 100,000 images in the datasets). From Tab. 1, we can observe that the ensembling still generally performs better than merging. The models on different datasets have different performances due to the dataset characteristics. Similarly to the model-wise comparison, with more models in the merging process (2 to 7 models merging results are in the supplementary materials), the better the performance for both merging and ensembling. The ensembling is more stable than merging in terms of the best and worst performance on different datasets. For example, on the PathMNIST [39, 40], uniform soup does not work very well (32.42% accuracy, which largely behind model ensemble methods) compared to its performance on the Cifar10/Cifar100 or Tiny ImageNet datasets. A similar situation shows on Celeba [17]: the uniformly average model only achieves 53.19%. Nevertheless, the greedy soup models can mitigate this problem (88.48% on PathMNIST and 93.06 on Celeba) as they can ensure a decent result by only considering the first model. These results may be caused by the contradictions between model weights.
@@ 6.2.3 Merging Models trained over different Tasks
Table 2: Performance Comparison for Task Merging Across Different Datasets with Various Model Averaging Methods. Model 1 is a VGG19 model trained on the Cifar100 with the classification accuracy of 70.37%, while Model 2 denotes the VGG19 model trained on different datasets, but evaluated on Cifar100 after merging/ensembling. E.g. Model 2 trained on Cifar10 receives 5.84% accuracy on Cifar100 dataset.
Datasets Data Size Model2 Uniform Soups Ens Lgt Ens Features
Cifar10 60,000 5.84 49.63 69.93 69.93
PathMNIST 107,180 1.29 24.25 69.88 70.00
DermaMNIST 10,015 5.77 60.37 69.06 69.36
Celeba 202,599 1.03 49.33 69.48 69.61
TinyImageNet 100,000 18.16 61.02 58.40 58.52
ChestXRay 112,120 15.18 61.38 68.18 68.33
We also explore model-merging between different well-trained models on different tasks, as shown in Tab. 2. Model 1 is a VGG19 model trained on the Cifar100 with the classification accuracy of 70.37%, while Model 2 denotes the VGG19 model trained on different datasets, but evaluated on Cifar100 after merging/ensembling. Generally, uniform souping within Tiny ImageNet, where Model 2 was trained on, has the highest accuracy of 61.02% that is better than the ensemble on logits (58.40% accuracy). It may be because the Tiny ImageNet dataset is a general and relatively large dataset, forcing Model 2 to learn a diverse set of features. Interestingly, Cifar10 should be a similar one to Cifar100, but the souping result does not outperform the model ensembling. Maybe because it is too small. In contrast, the ChestXRay dataset should be very different from Cifar100, but it shows a reasonably decent improvement with souping compared to Model 2 along. It illustrates that the dataset size also matters. The souping performance on Celeba (Model 2 is trained on) is not good, may be caused by the Celeba dataset only containing human faces that are very different from Cifar100 (Model 1 is trained on).